# These $\equalcontThese authors contributed equally to this work.$

### 27 Mar 2026 v1

@@ 1.1 Background
Let $(X,\omega)$ be a compact Hermitian manifold of complex dimension $n$, equipped with the Hermitian fundamental form $\omega$ given in local coordinates by $\omega=\frac{i}{2}\sum\limits_{\alpha,\beta}a_{\alpha,\bar{\beta}}dz_{\alpha}\wedge d\bar{z}_{\beta},$ where $(a_{\alpha,\bar{\beta}})$ is a positive definite Hermitian matrix. In the case $d\omega=0$, then $\omega$ is called a Kähler form and $(X,\omega)$ is said to be a compact Kähler manifold. The smooth volume form $dV$ associated to this Hermitian metric is given by the $n$-th wedge product $dV=\omega^{n}$. The study of the continuity of the complex Monge-Ampère operator $(dd^{c}\bullet)^{n}$ plays an important role in many problems of pluripotential theory. The original results in the direction of this research from the seminal work of Betford and Taylor in [1]. In this paper, Bedford and Taylor proved that if a sequence of locally bounded plurisubharmonic functions $(u_{j})\subset PSH(\Omega)\cap L^{\infty}_{loc}(\Omega)$, where $PSH(\Omega)$ denotes the set of plurisubharmonic functions on an open subset $\Omega\subset\mathbb{C}^{n}$ is decreasing to a locally bounded plurisubharmonic function $u$ then $(dd^{c}u_{j})^{n}$ is weakly convergent to $(dd^{c}u)^{n}$. They also established a result about weak convergence of the sequence of the complex Monge-Ampère $(dd^{c}u_{j})^{n}$ to $(dd^{c}u)^{n}$ under the assumption that $(u_{j})$ is an increasing sequence and a.e. convergent concerning the Lebesgue measure $dV_{2n}$ to a locally bounded plurisubharmonic $u$. In summary, by relying on the results obtained by Bedford-Taylor, we observe that the Monge-Ampère operator $(dd^{c}\bullet)^{n}$ is continuous in weak*-topology along monotonic sequences in the class of locally bounded plurisubharmonic functions.
@@
Note that, when studying the continuity of the Monge–Ampère operator $(dd^{c}u_{j})$ on Cegrell classes $\mathcal{E}(\Omega)$, it is usually necessary to assume that the sequence of functions $(u_{j})$ is bounded from below by a function $u_{0}$. However, when considering the continuity problem for the Monge–Ampère operator $(\omega+dd^{c}\cdot)$ on compact Kähler manifolds, as in the papers of Xing (refer [20, 21]) or Dinew and Hiep in ([8]), the authors were able to dispense with this condition, since finding such a lower bounding function is not always feasible. Now, regarding compact Hermitian manifolds, what results have been obtained in this direction of research? Unfortunately, results in this direction for the case $(X,\omega)$, which is a compact Hermitian manifold, remain quite limited up to now. Only recently one knew the study of weak convergence of the Monge-Ampère operator for the class of $\omega$-plurisubharmonic functions appeared, notably in the paper by Kołodziej and Ngoc Cuong Nguyen in [17]. Regrettably, in this paper, under the assumptions of convergence in capacity, the authors were only able to obtain weak convergence of a subsequence $(\omega^{n}_{u_{j_{k}}})$ of the original sequence $(\omega^{n}_{u_{j}})$ (see Theorem 1.1 in [17]). Inspired by the results in [17], in this paper, we aim to provide additional results on the weak convergence of the Monge-Ampère operator on compact Hermitian manifolds for the class of $\omega$-plurisubharmonic functions, in relation to convergence in $Cap_{\omega}$.
@@ 1.2 Organization
The paper is presented in four sections and organized as follows. In Section 2, we present important results concerning the class of $\omega$-plurisubharmonic functions ($\omega$-psh) that will be used in the subsequent parts of the paper. This includes the definition of the $\omega$-psh class, the Monge-Ampère operator for bounded $\omega$-psh functions, the notion of weak convergence of a sequence of Monge-Ampère operators, the definition of the capacity $Cap_{\omega}$ and the convergence of a sequence $(u_{j})$ of $\omega$-psh functions to an $\omega$-psh function in $Cap_{\omega}$, as well as several other related concepts. Section 3 is devoted to investigating the weak convergence of a sequence of the complex Monge-Ampère operator $(\omega+dd^{c}u_{j})^{n}$ to $(\omega+dd^{c}u)^{n}$ if we assume that $u_{j}$ is convergent to $u$ in $Cap_{\omega}$. In Section 4, we establish the existence of weak solutions of the Monge-Ampère equations on compact Hermitian manifolds under the assumption that there exists a smooth subsolution.

### 30 Mar 2026 v2

@@ 1.1 Background
Let $(X,\omega)$ be a compact Hermitian manifold of complex dimension $n$, equipped with the Hermitian fundamental form $\omega$ given in local coordinates by $\omega=\frac{i}{2}\sum\limits_{\alpha,\beta}a_{\alpha,\bar{\beta}}dz_{\alpha}\wedge d\bar{z}_{\beta},$ where $(a_{\alpha,\bar{\beta}})$ is a positive definite Hermitian matrix. In the case $d\omega=0$, then $\omega$ is called a Kähler form and $(X,\omega)$ is said to be a compact Kähler manifold. The smooth volume form $dV$ associated to this Hermitian metric is given by the $n$-th wedge product $dV=\omega^{n}$. The study of the continuity of the complex Monge-Ampère operator $(dd^{c}\bullet)^{n}$ plays an important role in many problems of pluripotential theory. The original results in the direction of this research from the seminal work of Betford and Taylor in [1]. In this paper, Bedford and Taylor proved that if a sequence of locally bounded plurisubharmonic functions $(u_{j})\subset PSH(\Omega)\cap L^{\infty}_{loc}(\Omega)$, where $PSH(\Omega)$ denotes the set of plurisubharmonic functions on an open subset $\Omega\subset\mathbb{C}^{n}$, is decreasing to a locally bounded plurisubharmonic function $u$ then $(dd^{c}u_{j})^{n}$ is weakly convergent to $(dd^{c}u)^{n}$. They also established a result about weak convergence of the sequence of the complex Monge-Ampère $(dd^{c}u_{j})^{n}$ to $(dd^{c}u)^{n}$ under the assumption that $(u_{j})$ is an increasing sequence and a.e. convergent concerning the Lebesgue measure $dV_{2n}$ to a locally bounded plurisubharmonic $u$. In summary, by relying on the results obtained by Bedford-Taylor, we observe that the Monge-Ampère operator $(dd^{c}\bullet)^{n}$ is continuous in weak*-topology along monotonic sequences in the class of locally bounded plurisubharmonic functions.
@@
Note that, when studying the continuity of the Monge–Ampère operator $(dd^{c}u_{j})^{n}$ on Cegrell classes $\mathcal{E}(\Omega)$, it is usually necessary to assume that the sequence of functions $(u_{j})$ is bounded from below by a function $u_{0}$. However, when considering the continuity problem for the Monge–Ampère operator $(\omega+dd^{c}\bullet)^{n}$ on compact Kähler manifolds, as in the papers of Xing (refer [20, 21]) or Dinew and Hiep in [8], the authors were able to dispense with this condition, since finding such a lower bounding function is not always feasible. Now, regarding compact Hermitian manifolds, what results have been obtained in this direction of research? Unfortunately, results in this direction for the case $(X,\omega)$, which is a compact Hermitian manifold, remain quite limited up to now. Only recently one knew the study of weak convergence of the Monge-Ampère operator for the class of $\omega$-plurisubharmonic functions appeared, notably in the paper by Kołodziej and Ngoc Cuong Nguyen in [17]. Regrettably, in this paper, under the assumptions of convergence in capacity, the authors were only able to obtain weak convergence of a subsequence $(\omega^{n}_{u_{j_{k}}})$ of the original sequence $(\omega^{n}_{u_{j}})$ (see Theorem 1.1 in [17]). Inspired by the results in [17], in this paper, we aim to provide additional results on the weak convergence of the Monge-Ampère operator on compact Hermitian manifolds for the class of $\omega$-plurisubharmonic functions, in relation to convergence in $Cap_{\omega}$.
@@ 1.2 Organization
The paper is presented in four sections and organized as follows.
In Section 2, we present important results concerning the class of $\omega$-plurisubharmonic functions ($\omega$-psh) that will be used in the subsequent parts of the paper. This includes the definition of the $\omega$-psh class, the Monge-Ampère operator for bounded $\omega$-psh functions, the notion of weak convergence of a sequence of Monge-Ampère operators, the definition of the capacity $Cap_{\omega}$ and the convergence of a sequence $(u_{j})$ of $\omega$-psh functions to an $\omega$-psh function in $Cap_{\omega}$, as well as several other related concepts. Section 3 is devoted to investigating the weak convergence of a sequence of the complex Monge-Ampère operator $(\omega+dd^{c}u_{j})^{n}$ to $(\omega+dd^{c}u)^{n}$ if we assume that $u_{j}$ is convergent to $u$ in $Cap_{\omega}$. In Section 4, we establish the existence of weak solutions of the Monge-Ampère equations on compact Hermitian manifolds under the assumption that there exists a smooth subsolution.